# October 17, 2025

### v2

@@ Abstract
We provide a complete description of unitarity bounds on the gauge-scalar sectors of models with extra $SU(2)$ doublet, neutral singlet, and charged singlet scalars. Such additions are very frequent in models beyond the Standard Model, and, in particular, they are almost universal in models explaining the dark matter problem. We propose a specific classification and minimal set of scattering matrices containing all the relevant information. We also developed a Mathematica implementation of our results, BounDS, and we use it to fully study a number of simple cases, comparing with the literature, when available. The Mathematica notebook BounDS is provided via a public GitHub repository.
@@ I Introduction
The Standard Model (SM) of particle physics has been extremely successful, culminating in the discovery of a scalar particle consistent with the Higgs boson ATLAS:2012yve ; CMS:2012qbp . Nevertheless, it is not a complete theory, and extensions beyond the Standard Model (BSM) are required. In particular, the SM provides no explanation for dark matter (DM) ParticleDataGroup:2016lqr , whose presence is inferred across a wide range of astronomical scales, from kiloparsecs to the size of the observable Universe. Key probes include the rotation curves of spiral galaxies, the study of galaxy clusters, and measurements of the Cosmic Microwave Background Planck:2018vyg ; Cirelli:2024ssz .
@@
One of the most widely studied scenarios envisions dark matter as consisting of new fundamental particles. The simplest scalar extension introduces a real, neutral $SU(2)$ singlet with a global $\mathbb{Z}_{2}$ symmetry to ensure stability Burgess:2000yq ; GAMBIT:2017gge . Alternatively, a minimal fermionic option is a Weyl fermion, commonly referred to as a sterile neutrino Cirelli:2024ssz . Other proposed models include the Inert Doublet Model, which extends the Higgs sector with a dark doublet Ilnicka:2015jba ; multi-singlet SM extensions Goncalves:2025snm ; the Next-to-Minimal Supersymmetric Standard Model Ellwanger:2009dp ; and models featuring scalar or fermionic portals to hidden sectors Arcadi:2024wwg . Most of these models include extra scalar fields, doublet or singlet under the SM gauge group, and often new discrete symmetries in order to stabilize the dark sector. In exploring such theories, one must start with the theoretical constraints, such as guaranty of a bounded from below potential, assurance of a global minimum, and perturbative unitarity.
@@
In any consistent quantum field theory, probability conservation is encoded in the unitarity of the $S$-matrix Schwartz:2014sze ; Logan:2022uus . In scalar extensions motivated by dark matter, this requirement becomes particularly relevant: new scalar fields can induce scattering amplitudes that grow with energy, and enforcing unitarity imposes bounds on their masses and couplings. For $2\to 2$ scattering processes, unitarity constraints can be applied to the partial-wave decomposition of the amplitudes, with the strongest high-energy limits coming from the zeroth partial-wave amplitude Schwartz:2014sze ; Logan:2022uus . In the high-energy limit, the Equivalence Theorem Veltman:1989ud ensures that the scattering of longitudinally polarized gauge bosons can be replaced by the corresponding Goldstone bosons, allowing these scalar amplitudes to capture the dominant contributions to unitarity bounds. Historically, Lee, Quigg, and Thacker applied this method to derive an upper bound on the Higgs boson mass, $M_{H}^{2}\leq 8\pi\sqrt{2}/(3G_{F})\approx 1\textrm{TeV}^{2}$ Lee:1977yc ; Lee:1977eg . Unitarity has also been applied to the general two-Higgs-doublet model (2HDM) with explicit CP violation Ginzburg:2005dt ; Kanemura:2015ska , to the general $N$-Higgs-doublet model Bento:2017eti , also with fermions Bento:2018fmy , and to constrain large scalar multiplets Hally:2012pu ; Milagre:2024wcg . More recently, perturbative unitarity analyses have been extended beyond the traditional $2\to 2$ limit. For instance, Ref. Bresciani:2025toe studies unitarity in general $M\to N$ scattering processes with $M,N\geq 2$. In particular, in inflationary models featuring a non-minimal coupling with the Ricci scalar, partial-wave unitarity bounds on $2\to N$ processes have improved earlier estimates of the energy scale at which new-physics effective interactions invalidate the perturbative expansion Steingasser:2025txd .
@@
In this work, we extend the analysis of Ref. Bento:2017eti by constructing an $SU(2)\times U(1)$-symmetric model with an arbitrary number of scalar fields (doublets, neutral singlets and charged singlets) and by imposing perturbative unitarity constraints in the high-energy limit. Using partial-wave unitarity for tree-level $2\to 2$ scattering, we derive bounds on the quartic couplings.
@@
In Section II, we present the most general renormalizable $SU(2)\times U(1)$-symmetric model in the high-energy limit, composed of an arbitrary number of $SU(2)$ scalar doublets with hypercharge $\tfrac{1}{2}$ and singlets that are either neutral or carry hypercharge 1. We derive constraints on the Lagrangian couplings imposed by hermiticity and identify additional symmetry relations arising from the interchange of field indices. In Section III, we construct the scattering matrix from second derivatives of the quartic potential, diagonalize it to obtain eigenamplitudes, and impose perturbative unitarity, $|\Lambda|\leq 8\pi$, to constrain the quartic couplings. We propose to classify states by the conserved quantum numbers $|Q,Y,T\rangle$, corresponding to electric charge, hypercharge, and total isospin, respectively. In Appendix A, we illustrate the benefits of labeling states not only by $Q$ and $Y$, but also by total isospin $T$, using the Standard Model as a case study in Appendix B. This classification facilitates the construction of scattering matrices and the identification of independent channels, for which we present a minimal basis (meaning that we discuss the selection of basis states that allows all independent eigenvalues to be determined). We provide explicit formulas for the potential and the matrix elements of all scattering processes allowed in our class of models. In Section IV, we present the Mathematica notebook BounDS that allows for the automatic calculation of the quartic part of the scalar potential and of all scattering matrices in models of this type; the user must only input the particle content, and, if needed, the transformation properties of the fields under additional flavour and/or generalized CP symmetries [^1] These can either be discrete or continuous, abelian or non-abelian. . We make the notebook publicly available at
@@
In Section V, we apply these methods to particular models by specifying the scalar content and symmetries, and we compare with results in the literature, when available. We present our results in terms of the parameters of the potential. In specific models, it is sometimes possible to write the parameters of the potential in terms of scalar masses and/or mixing angles. That involves deciding on terms of dimension two and/or three, defining the vacuum of the theory, writing all (neutral and charged) scalar mass matrices, diagonalizing them in order to obtain masses and mixing angles in terms of potential parameters, and then inverting those relations. This must be done on a case-by-case basis, and lies beyond the scope of this work. We present our conclusions in Section VI.
@@ II.1 Particle Content
Consider a gauge theory symmetric under the $SU(2)\times U(1)$ group with an arbitrary number of scalar fields (and their charge conjugates) that are either doublets or singlets of $SU(2)$. We define the electric charge $Q$ as the sum of the third component of isospin $T_{3}$, and the hypercharge $Y$, and restrict our discussion to models where the electric charge of the scalar particles is either singly charged [^2] In units of the electron charge $e$. or neutral. As a consequence, this theory may have $n_{D}$ $SU(2)$ scalar doublets with $Y=1/2$ which we denote by
$\Phi_{i}=\begin{pmatrix}\phi_{i}^{+},\,\phi_{i}^{0}\end{pmatrix}^{T},\quad i=1,\ldots,n_{D},$ (1)
@@
$\varphi_{i}^{+},\quad i=1,\ldots,n_{c},$ (2)
and $n_{n}$ $SU(2)$ real scalar singlets with $Y=0$ which we denote by
@@
Notice that models with $m$ $SU(2)$ complex scalar singlets with $Y=0$ are particular cases of models with $n_{n}=2m$ $SU(2)$ real scalar singlets with $Y=0$; hence, we do not treat them separately. The oblique radiative corrections for this class of models has been thoroughly studied in Refs. Grimus:2007if ; Grimus:2008nb .
@@ II.2 Scalar potential
In this work, we are interested in constraining the parameters of such models by imposing partial wave unitarity bounds on $2\to 2$ scattering processes. We perform all calculations in the high-energy limit where all contributions from propagators vanish, deeming the quartic interactions in the potential the dominant contributions. The most general renormalizable quartic part of the scalar potential of such a theory can be written as: [^3] For the simplest model including the $\kappa_{ab,cd}$ terms, see Sec. V.8 below.
@@
Although compact, not all couplings in this notation are independent. By swapping fields and requiring the scalar potential to be hermitian, we derive the following relations between quartic couplings:
@@ III.1 Partial wave decomposition
may have contributions from $s$, $t$, and $u$-channels, and contact interactions. However, as we take the limit where the Mandelstam variables $s$, $t$, and $u$ go to infinity, the contributions from $s$, $t$, and $u$-channels vanish, respectively  [^4] For an interesting example see, for example, Appendix A of MiguelMSc .. Consequently, in the high-energy limit, only the quartic interactions involving the external scalars contribute to the tree-level amplitude, allowing us to write it as
@@
Any square-integrable function may be expressed as an expansion in a complete set of basis functions. To study perturbative unitarity, it is particularly convenient to choose the basis of Legendre polynomials $P_{J}(\cos\theta)$, where $J$ is the total orbital angular momentum of the final state, and $\theta$ the scattering angle. Expanding the amplitude in this basis defines the so-called partial-wave expansion Hally:2012pu ; Milagre:2024wcg :
@@
and are constrained by tree-level partial-wave unitarity as Hally:2012pu ; Milagre:2024wcg
@@
Using the fact that, in the high-energy limit, the amplitude is independent of the scattering angle, and that $P_{0}(\cos\theta)=1$, the most stringent partial-wave unitarity bound arises by applying the second condition of Eq. (23) to the zeroth partial wave $a_{0}$. Thus, using Eq. (20), all quartic couplings must satisfy
@@ III.2 Scattering matrices
One may go a step further by employing the method of coupled-channel analysis Logan:2022uus ; Ginzburg:2005dt ; Kanemura:2015ska . This approach takes advantage of the fact that partial wave unitarity bounds may be imposed on any specific process, but also to any superposition of states, provided they have the same quantum numbers. By organizing the zeroth partial waves of such scatterings into a coupled-channel matrix, the unitarity condition translates into a bound on its eigenvalues. The most stringent constraint is obtained by requiring that the modulus of the largest eigenvalue to remain below $8\pi$. Once again, note that $J=0$ partial waves should be multiplied by a $1/\sqrt{2}$ for every pair of identical particles in the initial or final state.
We can classify two-particle initial and final states according to their total electric charge $Q$ and hypercharge $Y$, following the approach in Bento:2017eti . In addition to $Q$ and $Y$, any scattering process involving $SU(2)$ doublets must also conserve total isospin, $T$. Following the results of Appendix A, we list in Table 1 the most minimal set of independent two-particle states, labeled by the quantum numbers $|Q,Y,T\rangle$. Throughout we use the notation
@@
Looking at Table 1, we notice that the non-redundant sets of states could be classified exclusively in terms of $Y$ and $T$, as is done for the 2HDM in Ginzburg:2005dt . On the other hand, Ref. Bento:2017eti advocated $Q$ and $Y$ when considering all states, since then, as can be seen in Table 9 of Appendix A, there are states with different $Q$ for the same $(Y,T)$. We find it best to make the bridge and consider $|Q,Y,T\rangle$.
@@
$16\pi\,a_{0}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right$] $=$ $N_{ab}N_{cd}\,\left(2\lambda_{ca,db}+2\lambda_{da,cb}\right),$ (30)
$16\pi\,a_{0}\left[\phi^{+}_{[a}\,\phi^{0}_{b]}\to\phi^{+}_{[c}\,\phi^{0}_{d]}\right$] $=$ $\left(2\lambda_{ca,db}-2\lambda_{da,cb}\right),$ (31)
@@
Below, we study specific models, showing the scattering matrices $M_{|Q,Y,T\rangle}$ and, when an analytical expression is possible, the corresponding eigenvalues. Note that in Ref. Bento:2022vsb a method is proposed based on principal minors, that forgoes diagonalization. In specific numerical simulations, this is likely to be computationally preferable in all models dealing with large scattering matrices.

### v3

# Perturbative unitarity for models with singlet and doublet scalars

@@ Abstract
We provide a complete description of perturbative unitarity bounds on the gauge-scalar sectors of models with extra $SU(2)$ doublet, neutral singlet, and charged singlet scalars. Such additions are very frequent in models beyond the Standard Model, and, in particular, they are almost universal in models explaining the dark matter problem. We propose a specific classification and minimal set of scattering matrices containing all the relevant information. We also developed a Mathematica implementation of our results, BounDS, and we use it to fully study a number of simple cases, comparing with the literature, when available. The Mathematica notebook BounDS is provided via a public GitHub repository.
@@ I Introduction
The Standard Model (SM) of particle physics has been extremely successful, culminating in the discovery of a scalar particle consistent with the Higgs boson Aad and others (2012); Chatrchyan and others (2012). Nevertheless, it is unlikely to be the final theory, and extensions beyond the Standard Model are widely considered. In particular, the SM provides no explanation for dark matter Navas and others (2024), whose presence is inferred across a wide range of astronomical scales, from kiloparsecs to the size of the observable Universe. Key probes include the rotation curves of spiral galaxies, the study of galaxy clusters, and measurements of the Cosmic Microwave Background Aghanim and others (2020); Cirelli et al. (2024).
@@
One of the most widely studied scenarios envisions dark matter as consisting of new fundamental particles. The simplest scalar extension introduces a real, neutral $SU(2)$ singlet with a global $\mathbb{Z}_{2}$ symmetry to ensure stability Burgess et al. (2001); Athron and others (2017). Alternatively, a minimal fermionic option is a Weyl fermion, commonly referred to as a sterile neutrino Cirelli et al. (2024). Other proposed models include the Inert Doublet Model, which extends the Higgs sector with a dark doublet Ilnicka et al. (2016); multi-singlet SM extensions Gonçalves et al. (2025); the Next-to-Minimal Supersymmetric Standard Model Ellwanger et al. (2010); and models featuring scalar or fermionic portals to hidden sectors Arcadi et al. (2024). Most of these models include extra scalar fields, doublet or singlet under the SM gauge group, and often new discrete symmetries in order to stabilize the dark sector. Any consistent exploration of these theories must begin with the enforcement of basic theoretical requirements, such as the boundedness from below of the scalar potential, the existence of a global minimum, and perturbative unitarity, the latter being the main focus of the present work.
@@
In a consistent quantum field theory, probability conservation is encoded in the unitarity of the $S$-matrix Schwartz (2014); Logan (2022). In scalar extensions of the Standard Model, this requirement is particularly restrictive, as the presence of additional scalar degrees of freedom can lead to scattering amplitudes that grow with energy. Imposing perturbative unitarity, therefore, yields non-trivial constraints on the masses and couplings of the new scalars, ensuring the validity of perturbation theory. For $2\to 2$ scattering processes, perturbative unitarity constraints can be applied to the partial-wave decomposition of the amplitudes, with the strongest high-energy limits coming from the zeroth partial-wave amplitude Schwartz (2014); Logan (2022). In the high-energy limit, the equivalence theorem Veltman (1990) ensures that the scattering of longitudinally polarized gauge bosons can be replaced by the corresponding Goldstone bosons, allowing these scalar amplitudes to capture the dominant contributions to perturbative unitarity bounds. Historically, Lee, Quigg, and Thacker applied this method to derive an upper bound on the Higgs boson mass, $M_{H}^{2}\leq 8\pi\sqrt{2}/(3G_{F})\approx 1\textrm{TeV}^{2}$ Lee et al. (1977a, b). Perturbative unitarity has also been applied to the general two-Higgs-doublet model (2HDM) with explicit CP violation Ginzburg and Ivanov (2005); Kanemura and Yagyu (2015), to the general $N$-Higgs-doublet model Bento et al. (2017), also with fermions Bento et al. (2018), and to constrain large scalar multiplets Hally et al. (2012); Milagre and Lavoura (2024). Many other examples can be found in the literature, including Llewellyn Smith (1973); Cornwall et al. (1974); Weldon (1984); Casalbuoni et al. (1986, 1988); Gunion et al. (1991); Maalampi et al. (1991); Kanemura et al. (1993); Akeroyd et al. (2000); Horejsi and Kladiva (2006); El Kaffas et al. (2007); Grinstein et al. (2014); Nagai et al. (2015); Englert et al. (2016); Cacchio et al. (2016); Harris and Logan (2017); Goodsell and Staub (2018, 2019); Bento et al. (2023); Urquía-Calderón et al. (2025); Blažek et al. (2025). More recently, perturbative unitarity analyses have been extended beyond the traditional $2\to 2$ limit. For instance, Ref. Bresciani et al. (2025) studies unitarity bounds in general $M\to N$ scattering processes with $M,N\geq 2$. In particular, in inflationary models featuring a non-minimal coupling with the Ricci scalar, partial-wave unitarity bounds on $2\to N$ processes have improved earlier estimates of the energy scale at which new-physics effective interactions invalidate the perturbative expansion Steingasser et al. (2025).
@@
In this work, we extend the analysis of Ref. Bento et al. (2017) by constructing an $SU(2)\times U(1)$-symmetric model with an arbitrary number of scalar fields which may either be singlets or doublets under $SU(2)$. The hypercharge assignments are chosen such that the new scalars are either electrically neutral or carry a single unit of electric charge. With this framework, we then require all relevant $2\to 2$ scattering processes to satisfy perturbative unitarity at tree-level and in the high-energy limit, thereby deriving upper bounds on the quartic couplings of this class of theories.
@@
In Section II, we present the most general renormalizable $SU(2)\times U(1)$-symmetric model in the high-energy limit, composed of an arbitrary number of $SU(2)$ scalar doublets with hypercharge $\tfrac{1}{2}$ and singlets that may either be electrically neutral, or carry hypercharge 1. We derive relations among the couplings by imposing Hermiticity of the scalar potential and by eliminating redundant gauge-invariant operators. In Section III, we construct the scattering matrix from second derivatives of the quartic potential, diagonalize it to obtain eigenamplitudes, and impose perturbative unitarity, $|\Lambda|\leq 8\pi$, to constrain the quartic couplings. We propose to classify states by the conserved quantum numbers $|Q,Y,T\rangle$, corresponding to electric charge, hypercharge, and total isospin, respectively. In Appendix A, we illustrate the benefits of labeling states not only by $Q$ and $Y$, but also by total isospin $T$, using the Standard Model as a case study in Appendix B. This classification facilitates the construction of scattering matrices and the identification of independent channels, for which we present a minimal basis (meaning that we discuss the selection of basis states that allow all independent eigenvalues to be determined). We provide explicit formulas for the potential and the matrix elements of all scattering processes allowed in our class of models. In Section IV, we present the Mathematica notebook BounDS that allows for the automatic calculation of the quartic part of the scalar potential and of all scattering matrices in models of this type; the user must only input the particle content, and, if needed, the transformation properties of the fields under additional flavour and/or generalized CP symmetries [^1] These can either be discrete or continuous, Abelian or non-Abelian.. We make the notebook publicly available at
@@
In Section V, we apply these methods to particular models by specifying the scalar content and symmetries, and we compare with results in the literature, when available. We present our results in terms of the parameters of the potential. In specific models, it is sometimes possible to write the parameters of the potential in terms of scalar masses and/or mixing angles. That implies defining gauge-invariant terms of dimension two and/or three, defining the vacuum of the theory, writing all (neutral and charged) scalar mass matrices, diagonalizing them to obtain masses and mixing angles in terms of potential parameters, and then inverting those relations. This must be done on a case-by-case basis and lies beyond the scope of this work. We present our conclusions in Section VI.
@@ II.1 Particle Content
Consider a gauge theory symmetric under the $SU(2)\times U(1)$ group with an arbitrary number of scalar fields (and their charge conjugates) that are either doublets or singlets of $SU(2)$. We define the electric charge $Q$ as the sum of the third component of isospin $T_{3}$, and the hypercharge $Y$, and restrict our discussion to models where the scalar fields are either electrically neutral or carry a single unit of electric charge. As a consequence, this theory may have $n_{D}$ $SU(2)$ scalar doublets with $Y=1/2$ which we denote by
$\Phi_{i}=\begin{pmatrix}\phi_{i}^{+},\,\phi_{i}^{0}\end{pmatrix}^{T},\quad i=1,\ldots,n_{D};$ (1)
@@
$\varphi_{i}^{+},\quad i=1,\ldots,n_{c};$ (2)
$n_{n}$ $SU(2)$ real scalar singlets with $Y=0$ which we denote by
@@
Notice that models with $m$ $SU(2)$ complex scalar singlets with $Y=0$ are particular cases of models with $n_{n}=2m$ real scalar singlets with additional $\mathbb{Z}_{2}$ symmetries; hence, we do not treat them separately. The oblique radiative corrections for this class of models has been thoroughly studied in Refs. Grimus et al. (2008a, b).
@@ II.2 Scalar potential
In this work, we are interested in constraining the parameters of models like the one presented in Section II by imposing partial-wave unitarity bounds on $2\to 2$ scattering processes. We perform all calculations in the high-energy regime, where the contributions from propagators in the $s$, $t$, and $u$-channels vanish, deeming the quartic interactions in the potential the dominant contributions. The most general renormalizable quartic part of the scalar potential of such a theory can be written as: [^2] For the simplest model including the $\kappa_{ab,cd}$ terms, see Sec. V.8 below.
@@
Although compact, not all couplings in this notation are independent. By requiring the scalar potential to be Hermitian and rearranging gauge invariant bilinears [^3] By rearranging gauge invariant bilinears, we mean that $\left(\Phi_{a}^{\dagger}\Phi_{b}\right)\left(\Phi_{c}^{\dagger}\Phi_{d}\right)=\left(\Phi_{c}^{\dagger}\Phi_{d}\right)\left(\Phi_{a}^{\dagger}\Phi_{b}\right)\quad\Rightarrow\quad\lambda_{ab,cd}=\lambda_{cd,ab}.$ (5) , we derive the following relations between quartic couplings:
@@ III.1 Partial wave decomposition
may have contributions from $s$, $t$, and $u$-channels, and contact interactions. However, as we take the limit where the Mandelstam variables $s$, $t$, and $u$ go to infinity, the contributions from $s$, $t$, and $u$-channels vanish, respectively  [^4] For an interesting example, see, for instance, Appendix A of Bento (2017).. Consequently, in the high-energy limit, only the quartic interactions involving the external scalars contribute to the tree-level amplitude, allowing us to identify
@@
Any square-integrable function may be expressed as an expansion in a complete set of basis functions. To study perturbative unitarity, it is particularly convenient to choose the basis of Legendre polynomials $P_{J}(\cos\theta)$, where $J$ is the total angular momentum of the final state, and $\theta$ the scattering angle. Expanding the amplitude in this basis defines the so-called partial-wave expansion Hally et al. (2012); Milagre and Lavoura (2024):
@@
and are constrained by tree-level partial-wave unitarity through Hally et al. (2012); Milagre and Lavoura (2024)
@@
At tree-level, all $a_{J}$ are real, so the three inequalities in Eq. (24) reduce to $\left|a_{J}\right|\leq\frac{1}{2}$. In the high-energy limit, the scattering amplitudes are independent of the scattering angle, and since $P_{0}(\cos\theta)=1$, the strongest bound arises from the zeroth partial wave, $a_{0}$. The perturbative unitarity condition can then be written at the level of the amplitude as
$\left|a_{0}\right|\ =\ \frac{1}{16\pi}\Big|\mathcal{M}\left[A_{a}\,B_{b}\to C_{c}\,D_{d}\right]\Big|\ \leq\ \frac{1}{2},$ (25)
or, using Eq. (21), at the level of the quartic couplings as
@@ III.2 Scattering matrices
One may go a step further by employing the method of coupled-channel analysis Logan (2022); Ginzburg and Ivanov (2005); Kanemura and Yagyu (2015). This approach takes advantage of the fact that partial-wave unitarity bounds may be imposed on any specific process, but also to any superposition of states, provided they have the same quantum numbers. By organizing the zeroth partial waves of such scatterings into a coupled-channel matrix, the perturbative unitarity condition translates into a bound on its eigenvalues. The most stringent constraint is obtained by requiring that the modulus of the largest eigenvalue to remain below $8\pi$, as derived in Eq. (26). Once again, note that $J=0$ partial waves should be multiplied by a $1/\sqrt{2}$ for every pair of identical particles in the initial or final state.
We can classify two-particle initial and final states according to their total electric charge $Q$ and hypercharge $Y$, following the approach in Bento et al. (2017). In addition to $Q$ and $Y$, any scattering process involving $SU(2)$ doublets must also conserve total isospin, $T$. Following the results derived in Appendix A, we list in Table 1 the minimal set of independent two-particle states, labeled by the quantum numbers $|Q,Y,T\rangle$. Throughout we use the notation
@@
Looking at Table 1, we notice that the non-redundant sets of states could be classified exclusively in terms of $Y$ and $T$, as is done for the 2HDM in Ginzburg and Ivanov (2005). On the other hand, Ref. Bento et al. (2017) advocated for $Q$ and $Y$ when considering all states, since, as can be seen in Table 9 of Appendix A, there are states with different $Q$ for the same $(Y,T)$. Nevertheless, when considering scalars in representations of $SU(2)\times U(1)$ other than the ones we use here, the labeling of two-particle states by $Q$, $Y$, and $T$ is the most convenient as it leads to the minimal set of scattering matrices. Therefore, we find it best to make the bridge and consider $|Q,Y,T\rangle$.
@@
$16\pi\,a_{0}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right$] $=$ $2N_{ab}N_{cd}\,\left(\lambda_{ca,db}+\lambda_{da,cb}\right),$ (32)
$16\pi\,a_{0}\left[\phi^{+}_{[a}\,\phi^{0}_{b]}\to\phi^{+}_{[c}\,\phi^{0}_{d]}\right$] $=$ $2\left(\lambda_{ca,db}-\lambda_{da,cb}\right),$ (33)
@@
Below, we study specific models, showing the scattering matrices $M_{|Q,Y,T\rangle}$ and, when an analytical expression is possible, the corresponding eigenvalues. Note that in Ref. Bento et al. (2022) a method is proposed based on principal minors, that forgoes diagonalization. In specific numerical simulations, this is likely to be computationally preferable in all models dealing with large scattering matrices.